OpenStudy (anonymous):

how to simplify the equation sqrt(1+(f'(x))^2) to remove the sqrt when f(x)= ((3x^4)/8 )+ (1/(12x^2))

OpenStudy (anonymous):

some sort of "find the length of the curve" question?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

they really work hard to cook these questions up to get something you can find the anti-deritive of when you do the algebra

OpenStudy (anonymous):

i tried it once, it took me like half an hour to do it, maybe longer

OpenStudy (anonymous):

first you need the derivative did you get that?

OpenStudy (anonymous):

im not sure if im on the right track i am trying to leave it in terms of f'(x) until i can get rid of the sqrt

OpenStudy (anonymous):

i got 1^(1/2) + f'(x)^(3/2)

OpenStudy (anonymous):

you need to do it step by step a) find the derivative

OpenStudy (anonymous):

im really lost on this

OpenStudy (anonymous):

yeah i see

OpenStudy (anonymous):

lets go slow, there are no shortcuts

OpenStudy (anonymous):

thanks, you are a life saver

OpenStudy (anonymous):

before we begin, suppose you had to find \[\sqrt{1+6^2}\] what would you do?

OpenStudy (anonymous):

answer: you would a) square 6 b) add 1 c) take the square root that is what we have to do, no shortcuts

OpenStudy (anonymous):

a?

OpenStudy (anonymous):

first we need \(f'(x)\) i get \[f'(x)=\frac{9x^6-1}{6x^3}\]

OpenStudy (anonymous):

lol no, it wasn't multiple choice, you have to do all three in that order that is what we have to do a) take the derivative b) square it c) add 1 d) take the square root

OpenStudy (anonymous):

ah ok, mb

OpenStudy (anonymous):

so step one, as i said, take the derivative i cheated and used wolfram, but i am sure it is doable by hand if you have to "show your work"

OpenStudy (anonymous):

oh before i screw up totally, it is \[f(x)=\frac{3x^4}{8}+\frac{1}{12x^2}\] right ?

OpenStudy (anonymous):

yes it is

OpenStudy (anonymous):

ok then the derivative is easy enough

OpenStudy (anonymous):

its (9x^6 -1)/6x^3

OpenStudy (anonymous):

yes step two square it

OpenStudy (anonymous):

i might leave it as \[f'(x)^2=\frac{(9x^6-1)^2}{36x^6}\]

OpenStudy (anonymous):

yeah i just got that

OpenStudy (anonymous):

step 3 add 1

OpenStudy (anonymous):

sorry working on limited desk space

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

\[1+(f')^2=1+\frac{(9x^6-1)^2}{36x^6}=\frac{(9x^6-1)^2+36x^6}{36x^6}\]

OpenStudy (anonymous):

now the miracle occurs ready?

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

turns out this sucker is a perfect square ! (really imagine making this up)

OpenStudy (anonymous):

we can remove the denominator and the 36x^6 on the numerator with 1 right?

OpenStudy (anonymous):

if you see it, good, if not we can do the algebra

OpenStudy (anonymous):

oh no

OpenStudy (anonymous):

don't break it apart whatever you do !!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

lets do a tiny bit of algebra in the numerator

OpenStudy (anonymous):

\[(9x^6-1)^2+36x^6=81x^{12}-18x^6+1+36x^6=81x^{12}+18x^6+1=(9x^6+1)^2\] a total miracle

OpenStudy (anonymous):

oops that kind of bled over, hold on

OpenStudy (anonymous):

\[(9x^6-1)^2+36x^6\\=81x^{12}-18x^6+1+36x^6\\=81x^{12}+18x^6+1\\=(9x^6+1)^2\]

OpenStudy (anonymous):

in other words, this problem was cooked up so that the expression inside the radical is a perfect squre that isnot easyh to do

OpenStudy (anonymous):

you get \[\sqrt{\frac{(9x^6+1)^2}{36x^6}}\] take the square root now is easy

OpenStudy (anonymous):

let me know if you got stuck at the algebra

OpenStudy (anonymous):

1/6 (9 x^3+1/x^3) right?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

\[\frac{9x^6+1}{6x^3}\]

OpenStudy (anonymous):

kk

OpenStudy (anonymous):

don't take the square root of \(9x^6\) take the square root of \((9x^6+1)^2\) which is just \(9x^6+1\)

OpenStudy (anonymous):

i plugged it into wolfram and it gave me the 1/6 solution

OpenStudy (anonymous):

wasnt sure if was 9x^6+1 / 6x^3

OpenStudy (anonymous):

that is not the problem you can put the \(\frac{1}{6}\) out front if you like the problem you had was you changed the exponent on the \(9x^6\) term

OpenStudy (anonymous):

ok thanks this was very helpful

OpenStudy (anonymous):

that was wrong

OpenStudy (anonymous):

but whatever, it is easy enough now

OpenStudy (freckles):

I think ghost meant \[\frac{1}{6}(9x^3+\frac{1}{x^3}) \text{ which is the same as } \frac{9x^6+1}{6x^3}\]

OpenStudy (anonymous):

ooh

OpenStudy (anonymous):

the ghost of wolfram past

OpenStudy (anonymous):

yeah, wolfram tends to just confuse me more

OpenStudy (anonymous):

really it is amazing how they make these up take a derivative square it add 1 and voila, a perfect square !

OpenStudy (anonymous):

bet the lazy math teacher copies them out of the boook

OpenStudy (freckles):

in the calculus books I used to look up they would always put some kind of hint that 1+(f')^2 was a complete square

OpenStudy (anonymous):

probably

OpenStudy (anonymous):

@freckles go ahead an make one up without cheating i will time you lol

OpenStudy (anonymous):

no we were left in the dark, we were told it was possible that 1+f'(x) could have not been a perfect square

OpenStudy (anonymous):

and honestly that was what i was afraid of

OpenStudy (freckles):

lol \[f=\ln(\sec(x)) \\ f'=\frac{\sec(x) \tan(x)}{\sec(x)}=\tan(x) \\ 1+(f')^2=\sec^2(x)\]

OpenStudy (anonymous):

ok ok

OpenStudy (freckles):

but the ones that involve polynomials are harder to come up with

OpenStudy (freckles):

or rational expressions

OpenStudy (anonymous):

then you have to integrate secant i think i came up with \[f(x)=6\sqrt[3]{x^2}\] or something simple

OpenStudy (anonymous):

maybe that doesn't work dang

OpenStudy (anonymous):

another reason why i loathe calc 2

OpenStudy (freckles):

I think also a good thing to remember is that \[(f-\frac{1}{4f})^2 =f^2-2 \cdot f \frac{1}{2f}+(\frac{1}{4f})^2 \\ (f-\frac{1}{4 f})^2=f^2-\frac{1}{2}+(\frac{1}{4f})^2 \\ \\ \text{ and doing 1 plus on both sides gives } \\ (f+\frac{1}{4f})^2=f^2+\frac{1}{2}+(\frac{1}{4f})^2\] like this is what happened on this one when expanded the (f')^2 thing you had a -1/2 for the middle term \[f(x)=\frac{3}{8}x^4+\frac{1}{12 x^2 } \\ f'(x)=\frac{3}{2}x^3-\frac{1}{6 x^3} \\ (f')^2=(\frac{3}{2}x^3-\frac{1}{6x^3})^2 \text{ expanding this would give middle term is } \frac{-1}{2} \\ \text{ so adding one to the } (f')^2 \text{ will just mean you have } \\ \text{ to change the middle sign } \\ (f')^2+1=(\frac{3}{2}x^3+\frac{1}{6x^3})^2\]

OpenStudy (freckles):

oops I made a type in my first code line

OpenStudy (freckles):

that middle should be -2*f*1/(4 f) but everything that seems good and maybe I shouldn't have used f

OpenStudy (anonymous):

like the last one nice

OpenStudy (anonymous):

doesn't mean i don't still hate calc 2 though

OpenStudy (freckles):

So to make a problem... \[\text{ we want } f'(x)=a x^{n}-\frac{1}{b x^{n}} \\ \text{ where } \frac{a}{b}=\frac{1}{4} \\ (f'(x))^2=(ax^n)^2-2 ax^n \frac{1}{b x^n}+(\frac{1}{b x^n})^2 \\ (f'(x))^2=(ax^n)^2-\frac{1}{2}+(\frac{1}{bx^n})^2 \\ \text{ so } 1+(f'(x))^2=(ax^n)^2+\frac{1}{2}+(\frac{1}{b x^n})^2 =(ax^n+\frac{1}{b x^n})^2\] so integrating f' we can find an f to use

OpenStudy (freckles):

\[f'(x)=ax^n-\frac{1}{b x^n} \text{ where } \frac{a}{b}=\frac{1}{4} \\ f'(x)=ax^n-\frac{1}{b}x^{-n} \\ f(x)=a \frac{x^{n+1}}{n+1}- \frac{1}{b} \frac{x^{-n+1}}{-n+1}+C \\ \text{ where } n \neq -1,1\]

OpenStudy (anonymous):

ok show off!

OpenStudy (anonymous):

guess my challenge was more than met

OpenStudy (freckles):

lol :p I bet there could be other choices though for f (besides trigonometric like the one I chose earlier)